# Coarsening Kinetics in Active Model B+: Macroscale and Microscale Phase Separation

### v1

@@ Abstract
We perform a comprehensive numerical investigation of the coarsening kinetics of active Brownian particles modeled by the Active Model B+ (AMB+). This model was introduced by Tjhung et al. [Phys. Rev. X 8, 031080 (2018)] and is a generalization of Model B for a conserved order parameter, with two additional activity terms. These terms correspond to rotation-free current (of strength $\lambda$) and rotational current (of strength $\xi$). We find that the presence of rotational current $(\xi\neq 0)$ significantly affects growth kinetics. Depending on the parameter values, AMB+ exhibits either macroscale phase separation (MPS) or microscale phase separation ($\mu$PS). We present detailed results for the kinetics of MPS and $\mu$PS in AMB+ with critical composition.
@@ 1 Introduction
Another intriguing class of systems consists of active colloids that lack orientational order, such as motile colloids with spherical shapes (e.g. synthetic Janus particles). The alignment interactions in these systems are usually negligible, and the particles can be considered isotropic. As a result, these particles with purely repulsive interactions undergo liquid-gas phase separation, which is not possible for passive colloids without attractive forces. This process is known as MIPS and is the result of a nonequilibrium mechanism due to intrinsic motility Stenhammar et al. (2013); Cates and Tjhung (2018); Caporusso et al. (2020); Mandal et al. (2019); Caporusso et al. (2020); Shi et al. (2020); Martin-Roca et al. (2021). For example, bacteria with quorum sensing and active Janus particles with density-dependent self-propulsion speed exhibit MIPS Worlitzer et al. (2021); Sokolov et al. (2007). The underlying mechanism is due to the lack of microscopic time-reversal symmetry (TRS) in these systems. Consequently, the steady state of such systems does not obey the principle of detailed balance.
@@ 2 Model and Simulation Details
Coarsening kinetics in passive systems (e.g. binary mixtures) is described by a dynamical field theory with a conserved order parameter field, $\psi({\bf r},t)$, where ${\bf r}$ denotes space. Here, $\psi$ represents the local concentration difference between the two species. These systems obey detailed balance and preserve TRS. For passive systems with diffusive transport, the evolution of $\psi$ obeys the continuity equation or MB Hohenberg and Halperin (1977):
@@
[…]
In Eq. (3), $a$, $b$, and $K$ are positive constants, and $T$ is the temperature. The critical temperature of phase separation is denoted as $T_{c}$. The first two terms in the RHS represent the bulk free energy density. The third term represents the energy cost (surface tension) associated with variations in the order parameter. MB governs the evolution of $\psi$ to the free energy minima
$\psi_{0}=\pm\left[\frac{a}{b}(T_{c}-T)\right]^{1/2},$ (4)
which correspond to the equilibrium phases for $T<T_{c}$.
@@
$-\psi_{s}+\psi_{s}^{3}-\frac{d^{2}\psi_{s}}{dz^{2}}+\left(\lambda-\frac{\xi}{2}\right)\left(\frac{d\psi_{s}}{dz}\right)^{2}=\mu_{s},$ (9)
@@
$-\psi_{s}+\psi_{s}^{3}=\mu_{s}(\alpha).$ (11)
It is straightforward (though messy) to solve this cubic equation. The corresponding perturbative result is
@@
[…]
@@ 3.1 Macroscale Phase Separation (MPS)
[…]
@@
where $f(x)$ is a scaling function. We define ${L(t)}$ as the distance at which ${C(r,t)}$ first decays to zero. A crucial quantity to understand the asymptotic growth law is the effective dynamic exponent, defined as
@@
To quantify the asymptotic growth law, we calculate ${z_{\mathrm{eff}}}$ as a function of $t$. In Figs. 2(b)-2(e), we plot $1/z_{\mathrm{eff}}$ vs. $t$. For $\xi=0$, $z_{\mathrm{eff}}$ shows a flat trend with fluctuations around the expected value $31/z_{\mathrm{eff}}=1/$ 3. We should remark that the data for $z_{\rm eff}$ is usually noisy due to the numerical approximation of the derivative in Eq. (15). For $\xi<0$, at early times, all plots are flat around $31/z_{\mathrm{eff}}=1/$ 3, followed by a dip to values smaller than $1/4$ for $t>t_{c}(\xi)$. At late times, all plots tend towards the exponent $1/4$. A couple of remarks are in order here. First, the crossover time $t_{c}$ has an inverse dependence on $\xi$. We define $t_{c}$ as the time where $1/z_{\rm eff}$ first falls to $1/4$ in Fig. 2(c)-(e), and plot $t_{c}$ vs. $\xi$ on a log-log scale in Fig. 3. The data show an approximate power law behavior with $t_{c}\sim\xi^{-2}$, although the range of $\xi$ is limited. Our second remark is that AMB+ shows $t^{1/3}$-growth in the window where the morphologies are bicontinuous in Fig. 1, whereas the system shows $t^{1/4}$-growth in the asymptotic regime where the morphology is droplet-like.
@@
In Fig. 5, we study the direction of the current vector field for $\xi=-2$ at $t=10^{5}$. The direction of the vector ($\theta\in[-\pi,\pi]$) is shown using a colormap in Fig. 5(a), where only vectors with $J\geq aJ_{m}$ are shown for clarity ($a=0.02$, $J_{m}=0.022$ is the maximum current magnitude). As expected, the bulk current is negligible, whereas interfacial currents form localized closed loops. A zoomed-in view around a single droplet (marked X) in Fig. 5(b) highlights these circulating flows. To quantify this, Fig. 5(c) shows the curl of the current, revealing alternating clockwise and anticlockwise current loops along the interface. These surface loops redistribute the mass tangentially. For droplets of varying sizes, the loops carry currents of different magnitudes. Material is transported from smaller droplets to larger droplets via these unbalanced surface loops, thereby driving forward Ostwald ripening. The total number of vortices is even due to the periodic domain, which enforces zero net vorticity, consistent with Stokes’ theorem under periodic boundary conditions.
In MB, ripening is driven by the evaporation of small droplets, diffusive transport through the surrounding region, and condensation on larger droplets. This evaporation-condensation mechanism yields the LS growth law. Thus, tiny currents in bulk domains drive growth. In AMB+, bulk currents are swamped by interfacial currents that drive growth through surface diffusion. The corresponding growth mechanism has been clarified by Puri et al. Puri et al. (1997) and van Gemmert et al. Van Gemmert et al. (2005), and yields $L(t)\sim t^{1/4}$.
@@
[…]
@@
3.2 Microscale Phase Separation ($\mu$PS)
Next, we present numerical results for $\lambda<0$ and $\xi<0$. Our parameter choices lie in the region labeled reverse Ostwald ripening in the phase diagram of Tjhung et al. (2018). In Fig. 9 (top frames), we show steady-state snapshots at $t=10^{4}$ for $\lambda=-4$ and $\xi=-1$, $-2$. As in Fig. 1, the color bar indicates the range of $\psi$. As $\alpha=\lambda-\xi/2<0$ in this case, $\psi_{1}<|\psi_{2}|$ in Eq. (12), so the $\psi<0$ phase is the “minority phase”. However, the snapshots show static droplets of $\psi_{1}$ immersed in a background of $\psi_{2}$. Thus, the droplet morphology in Fig. 9 (top frames) is not an off-critical morphology in the sense of MB. Rather, the droplets occupy a larger fraction of the system than the background.
@@ 3.2 Microscale Phase Separation ($\mu$PS)
Can we make some quantitative statements about the hexagonal morphologies in the upper frames of Fig. 9? First, due to the conservation constraint, we expect (for $d=2$)
$3\pi L_{s}^{2}\psi_{1}+\left(\frac{3\sqrt{3}a^{2}}{2}-3\pi L_{s}^{2}\right)\psi_{2}$ $=$ $0,$ (19)
where $a$ is one of the sides of the hexagonal cell. Thus, we have
@@
Second, what is the dependence of $L_{s}(\lambda,\xi)?$ For small values of $\lambda$ or $\xi$, $L_{s}$ becomes very large. Thus, simulations on any reasonable lattice size encounter finite-size effects before we reach the saturation scale. On the other hand, for large values of $\lambda$ and $\xi$, simulations with the current mesh sizes become numerically unstable. This leaves a limited range of values for $\lambda$ and $\xi$ where we can drive the system to the saturated morphology. In Fig. 10(b), we plot $L_{s}$ vs. $-\xi$ on a log-log scale for $\lambda=-2.5,-3,-4$. The data appear to be consistent with the power law $L_{s}\sim(-\xi)^{-2/3}$ in the small $\xi$ range studied. In Fig. 10(c), we plot $L_{s}$ vs. $-\lambda$ for $\xi=-1.5,-1.8$ and $-2.0$. In this case, $L_{s}$ shows a more complicated non-monotonic behavior as $L_{s}\rightarrow\infty$ for $|\lambda|\rightarrow 0$. Simple scaling arguments based on a single droplet do not yield an understanding of the dependence of $L_{s}(\lambda,\xi)$. We believe that a reasonable interpretation of Figs. 10(b)-(c) requires an understanding of the multi-droplet interactions in Fig. 9.
In Fig. 11, we study the correlation function for the morphologies shown in Fig. 9. We plot $C(r,t)$ vs. $r/L$ at $t=10^{4}$ (in the saturated regime) for $\lambda=-2.5,-4$ and $\xi=-1,-2$. The scaled correlation function for $\mu$PS is independent of $\lambda$ and $\xi$, and is superuniversal. The corresponding snapshots show a hexagonal morphology for all these $(\lambda,\xi)$-values. The only difference is the magnitude of $L_{s}$. The scaling function shows marked oscillations, which reflect the crystalline nature of the $\mu$PS morphology.
@@ 4 Summary and Conclusion
AMB+ has a rich phenomenology and exhibits macro-phase separation (MPS) or micro-phase separation ($\mu$PS) in different regions of $(\lambda,\xi)$ space; see Fig. 4 in Ref. Tjhung et al. (2018). In the former case (MPS), the domain scale diverges with a power-law behavior: $L(t)\sim t^{\theta}$. Our numerical studies show that $\theta\simeq 1/4$ in the asymptotic regime. The current peaks sharply at the boundaries in the form of circulating loops, and the dominant growth mechanism is surface diffusion Van Gemmert et al. (2005); Puri et al. (1997). The presence of significant currents at the domain interfaces is found in other nonequilibrium systems as well, and is a signature of a system settling into a steady state Venkatareddy et al. (2025). In this parameter range, the system evolves by forward Ostwald ripening, with larger droplets growing at the cost of smaller droplets. Because of the distribution of droplet sizes, there is no cancellation of current loops at the droplet boundaries, leading to ongoing macroscopic transport and growth. In addition, evolution morphologies exhibit dynamical scaling, but the asymptotic scaling function depends on $\lambda$ and $\xi$, that is, it is not superuniversal. This is due to the asymmetry between the two phases. Although the overall composition is critical, the morphologies are not bicontinuous. This off-criticality results in the dependence of the scaling function on $\lambda$ and $\xi$. Nevertheless, there are several universal features, e.g., Porod’s law due to sharp interfaces.
For the case of $\mu$PS, the emergent morphologies are more subtle. In this case, the kinetics is driven by reverse Ostwald ripening, with smaller droplets growing at the expense of larger droplets. This equalizes the droplet sizes, with a resultant cancellation of circulating currents at the interfaces. In $d=2$, the coarsening system saturates to a morphology in which droplets of size $L_{s}(\lambda,\xi)$ are located at the vertices of a hexagonal lattice. The lattice spacing has a simple dependence on $L_{s}$, which is obtained from the conservation law. We believe that a complete understanding of the asymptotic $\mu$PS morphology must account for multi-droplet interactions. Simple arguments based on single-droplet growth are not adequate to understand the dependence of $L_{s}$ on $\lambda$ and $\xi$.

### v2

@@ Abstract
[…]
@@ 1 Introduction
Another intriguing class of systems consists of active colloids that lack orientational order, such as motile colloids with spherical shapes (e.g. synthetic Janus particles). Alignment interactions in these systems are usually negligible and particles can be considered geometrically isotropic, in contrast to elongated or rod-like active matter. These particles with purely repulsive interactions undergo liquid-gas phase separation, which is not possible for passive colloids without attractive forces. This process is known as MIPS and is the result of a nonequilibrium mechanism due to intrinsic motility Stenhammar et al. (2013); Cates and Tjhung (2018); Caporusso et al. (2020); Mandal et al. (2019); Caporusso et al. (2020); Shi et al. (2020); Martin-Roca et al. (2021). For example, bacteria with quorum sensing and active Janus particles with density-dependent self-propulsion speed exhibit MIPS Worlitzer et al. (2021); Sokolov et al. (2007). The underlying mechanism is due to the lack of microscopic time-reversal symmetry (TRS) in these systems. Consequently, the steady state of such systems does not obey the principle of detailed balance.
@@ 2 Model and Simulation Details
Coarsening kinetics in passive systems (e.g. binary mixtures without activity) is described by a dynamical field theory with a conserved order parameter field, $\psi({\bf r},t)$, where $\bf r$} denotes space. Here, $\psi$ represents the local concentration difference between the two species. These systems obey detailed balance and preserve TRS. For passive systems with diffusive transport, the evolution of $\psi$ obeys the continuity equation or MB Hohenberg and Halperin (1977):
@@
$\mathcal{F}[\psi({\bf r},t)]$ $=$ $\int{d{\bf r}\left\{\frac{a(T-T_{c})}{2}\psi^{2}+\frac{b}{4}\psi^{4}+\frac{K}{2}\left|{\nabla}\psi\right|^{2}\right\}}.$ (3)
In Eq. (3), $a$, $b$, and $K$ are positive constants, and $T$ is the temperature. The critical temperature of phase separation is denoted as $T_{c}$. The first two terms in the RHS represent the bulk free energy density. The third term represents the energy cost (surface tension) associated with variations in the order parameter. For $T<T_{c}$, MB governs the evolution of $\psi$ to the free energy minima
$\psi_{0}=\pm\left[\frac{a}{b}(T_{c}-T)\right]^{1/2}.$ (4)
For a general form of the free energy, the equilibrium phases are obtained from the well-known Maxwell common tangent construction, which ensures the same chemical potential and pressure between the coexisting phases.
@@
The phase diagram of Eq. (7) in the $(\lambda,\xi)$-plane is shown in Fig. 4 of Ref. Tjhung et al. (2018). There is a region of ($\lambda,\xi$)-values where the system settles into steady-state droplet configurations with monodisperse sizes. This region corresponds to reverse Ostwald ripening, where smaller droplets grow at the expense of larger ones. Here, the density outside the larger droplets exceeds that around the smaller droplets, causing mass flux from large to small. In contrast, there is another region of ($\lambda,\xi)$-values where the system undergoes forward Ostwald ripening, with smaller droplets growing at the expense of larger ones, eventually leading to bulk phase separation.
@@
$\psi_{s}+\psi_{s}^{3}-\frac{d^{2}\psi_{s}}{dz^{2}}+\left(\lambda-\frac{\xi}{2}\right)\left(\frac{d\psi_{s}}{dz}\right)^{2}=\mu_{s},$ (9)
@@
$\psi_{s}+\psi_{s}^{3}=\mu_{s}(\alpha).$ (11)
It is straightforward (though messy) to solve this cubic equation. The corresponding perturbative result is Wittkowski et al. (2014)
@@
The initial condition consists of uniformly distributed small-amplitude fluctuations about 0: $\psi({\bf r},0)=0+\delta\psi({\bf r},0)$, where $\delta\psi\in[-0.05,0.05]$. This mimics the disordered state with critical composition prior to the “quench”. The system is spontaneously unstable and undergoes phase separation via spinodal decomposition. The conservative nature of AMB+ ensures that $\int\psi({\bf r},t)~d{\bf r}=0$ for all $t$. All other cases with a nonzero average of $\psi$ are known as off-critical mixtures and correspond to asymmetric binary mixtures.
@@ 3.1 Macroscale Phase Separation (MPS)
where $\langle\ldots\rangle$ denotes an average over independent initial conditions. As our system is isotropic, $C({\bf r},t)$ can be spherically averaged to obtain $C(r,t)$. This quantity exhibits dynamical scaling if domain growth is characterized by a unique length scale L(t){L(t)} Puri and Wadhawan (eds.):
@@
where $f(x)$ is a scaling function. We define L(t){L(t)} as the distance at which C(r,t){C(r,t)} first decays to zero. A crucial quantity to understand the asymptotic growth law is the effective dynamic exponent, defined as
@@
To quantify the asymptotic growth law, we calculate $z_{\mathrm{eff}}$} as a function of $t$. In Figs. 2(b)-2(e), we plot $1/z_{\mathrm{eff}}$ vs. $t$. For $\xi=0$, $z_{\mathrm{eff}}$ shows a flat trend with fluctuations around the expected value $1/z_{\mathrm{eff}}=1/3$. We should remark that the data for $z_{\rm eff}$ is usually noisy due to the numerical approximation of the derivative in Eq. (15). For $\xi<0$, at early times, all plots are flat around $1/z_{\mathrm{eff}}=1/3$, followed by a dip to values smaller than $1/4$ for $t>t_{c}(\xi)$. At late times, all plots tend towards the exponent $1/4$. A couple of remarks are in order here. First, the crossover time $t_{c}$ has an inverse dependence on $\xi$. We define $t_{c}$ as the time where $1/z_{\rm eff}$ first falls to $1/4$ in Fig. 2(c)-(e), and plot $t_{c}$ vs. $\xi$ on a log-log scale in Fig. 3. The data show an approximate power law behavior with $t_{c}\sim\xi^{-2}$, although the range of $\xi$ is limited. Our second remark is that AMB+ shows $t^{1/3}$-growth in the window where the morphologies are bicontinuous in Fig. 1, whereas the system shows $t^{1/4}$-growth in the asymptotic regime where the morphology is droplet-like. A similar droplet morphology arises for MB with off-critical composition $\psi_{0}\neq 0$. However, in that case, the growth of the droplets continues to obey the LS law.
@@
In Fig. 5, we study the direction of the current vector field for $\xi=-2$ at $t=10^{5}$. The direction of the vector ($\theta\in[-\pi,\pi]$) is shown using a colormap in Fig. 5(a), where only vectors with $J\geq aJ_{m}$ are shown for clarity ($a=0.02$, $J_{m}=0.022$ is the maximum current magnitude). As expected, the bulk current is negligible, whereas interfacial currents form localized closed loops. A zoomed-in view around a single droplet (marked X) in Fig. 5(b) highlights these circulating flows. To quantify this further, Fig. 5(c) shows the curl of the current, revealing alternating clockwise and anticlockwise current loops along the interface. These loops are separated by nodes where the current is negligible. For droplets of varying sizes, the loops carry currents of different magnitudes. Material is transported from smaller droplets to larger droplets via these unbalanced surface loops, thereby driving forward Ostwald ripening. This is illustrated in Fig. 5(d), where we magnify the lower left-hand corner of Fig. 5(b). The arrows now denote the current at points where $J\geq 0.05J_{m}$, i.e., the cut-off is less stringent than in (b) where $J\geq 0.57J_{m}$. The total number of vortices is even due to the periodic domain, which enforces zero net vorticity, consistent with Stokes’ theorem under periodic boundary conditions.
In MB, ripening is driven by the evaporation of small droplets, diffusive transport through the surrounding region, and condensation on larger droplets. This evaporation-condensation mechanism yields the LS growth law. Thus, tiny currents in bulk domains drive growth. In AMB+, bulk currents are swamped by interfacial currents that drive growth through surface diffusion. In the process of mass transfer, the material has to move along a loop on the surface of a smaller droplet before joining a loop on the surface of a larger droplet. These loops become longer in time, scaling with the droplet size $L$. This tortuous process slows down domain growth as follows. The additional distance covered by the mass is the loop length $\sim L$, renormalizing the effective time to $\tau\sim t/L$, where we have assumed ballistic transport along the loops. For diffusive transport, we expect $L\sim\tau^{1/3}$, yielding the observed growth law $L\sim t^{1/4}$. The same growth law has been derived in the context of MB driven by surface diffusion Puri et al. (1997); Van Gemmert et al. (2005). In Ref. Puri et al. (1997), the role of surface diffusion is amplified by introducing an order-parameter dependent mobility $M(\psi)=1-\psi^{2}$ which sets the mobility in the bulk to zero. Such a physical situation arises, e.g., when the bulk phase undergoes gel or glass formation.
@@
In Fig. 6(a)-(b), we plot $C(r,t)$ vs. $r/L$ for $\xi=-2$. Figs. 6(a) and 6(b) show the scaling plot for the early-time regime $t<t_{c}$ $(t=10^{2},10^{3})$ and the late-time regime $t>t_{c}$ $(t=10^{4},10^{5},10^{6})$, respectively. There is a good data collapse in both regimes. Furthermore, the early-time scaling function is in excellent agreement with the MB function, denoted by a solid line in Fig. 6(a)-(b). However, subtle differences appear in the late-time scaling function of AMB+. For example, the oscillations (dictated by the conservation law $\int d{\bf r}~C({\bf r},t)=0$) have a smaller amplitude for AMB+. This is a consequence of the crossover in morphology. At first, the morphology of AMB+ is bicontinuous for all $\xi$ as well as MB. Therefore, the scaling is superuniversal, that is, independent of $\xi$. This is confirmed in Fig. 6(c), where we plot the early-time scaling function for $\xi=0$ (MB), 1,2,3-1,-2,-3. At late times, the bulk domains in AMB+ saturate to their preferred equilibrium values, which are asymmetric. Due to this asymmetry, the fractions of the two phases are $\phi_{1}=-\psi_{2}/(\psi_{1}-\psi_{2})$ and $\phi_{2}=\psi_{1}/(\psi_{1}-\psi_{2})$, respectively. Thus, the AMB+ morphology at late times is analogous to the MB morphology with an off-critical composition. For MB, it is well known that the scaling function changes continuously with the off-criticality Oono and Puri (1987), although its analytic form remains an outstanding problem. Thus, we expect the late-time scaling function of AMB+ to continuously vary with $\xi$. This is seen in Fig. 6(d), where we plot the asymptotic scaling function for $\xi=0,-1,-2,-3$.
@@
[…]
Next, we present numerical results for $\lambda<0$ and $\xi<0$. Our parameter choices lie in the region labeled reverse Ostwald ripening in the phase diagram of Tjhung et al. (2018). In Fig. 9 (top frames), we show steady-state snapshots at $t=10^{4}$ for $\lambda=-4$ and $\xi=-1$, 2-2. As in Fig. 1, the color bar indicates the range of $\psi$. As $\alpha=\lambda-\xi/2<0$ in this case, $\psi_{1}<|\psi_{2}|$ in Eq. (12), so the $\psi<0$ phase is the “minority phase”. However, the snapshots show static droplets of $\psi_{1}$ immersed in a background of $\psi_{2}$. Thus, the droplet morphology in Fig. 9 (top frames) is not an off-critical morphology in the sense of MB. Rather, the droplets occupy a larger fraction of the system than the background.
@@
Can we make some quantitative statements about the hexagonal morphologies in the upper frames of Fig. 9? First, consider a typical hexagonal unit cell of side $a$, whose vertices are located at the centers of saturated droplets arranged in a regular lattice. Within this cell, we compute the areas occupied by the two phases, $\psi_{1}$ and $\psi_{2}$. For a critical composition, the sum of the contributions from both regions must cancel. Therefore, we expect (for $d=2$)
$3\pi L_{s}^{2}\psi_{1}+\left(\frac{3\sqrt{3}a^{2}}{2}-3\pi L_{s}^{2}\right)\psi_{2}$ $=$ $0.$ (19)
Thus, we have
@@
Second, what is the dependence of $L_{s}(\lambda,\xi)?$ For small values of $\lambda$ or $\xi$, $L_{s}$ becomes very large. Thus, simulations on any reasonable lattice size encounter finite-size effects before we reach the saturation scale. On the other hand, for large values of $\lambda$ and $\xi$, simulations with the current mesh sizes become numerically unstable. This leaves a limited range of values for $\lambda$ and $\xi$ where we can drive the system to the saturated morphology. In Fig. 10(b), we plot $L_{s}$ vs. $\xi$ on a log-log scale for $\lambda=-2.5,-3,-4$. The data appear to be consistent with the power law $L_{s}\sim(-\xi)^{-2/3}$ in the small $\xi$ range studied. In Fig. 10(c), we plot $L_{s}$ vs. $\lambda$ for $\xi=-1.5,-1.8$ and 2.0-2.0. In this case, $L_{s}$ shows a more complicated non-monotonic behavior as $L_{s}\rightarrow\infty$ for $|\lambda|\rightarrow 0$. Simple scaling arguments based on a single droplet do not yield an understanding of the dependence of $L_{s}(\lambda,\xi)$. We believe that a reasonable interpretation of Figs. 10(b)-(c) requires an understanding of the multi-droplet interactions in Fig. 9.
[…]
@@ 4 Summary and Conclusion
[…]
For ease of reference, we summarize in Table 1 the domain morphologies and coarsening laws for the different models discussed in this paper, highlighting the distinct growth behaviors in AMB, AMB+, and their passive counterpart MB.
Model Morphology Growth Exponent Process
MB with MPS Bicontinuous $1/3$ Ostwald ripening
AMB with MPS Droplet $1/3\to 1/4$ Ostwald ripening
AMB+ with MPS Droplet $1/3\to 1/4$ Ostwald ripening
[…]
Table 1: Summary of morphologies and growth laws for different models with $\psi_{0}=0$.